OpenStudy (anonymous):

Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5

OpenStudy (anonymous):

i got c

OpenStudy (anonymous):

Yeah that what i thought too thx

OpenStudy (anonymous):

you welcome jus tag me if u need more help

OpenStudy (nthenic_oftime):

c :)

OpenStudy (anonymous):

What about this one Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

it looks the same im confused

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

What value of k makes the statement true? –(k + 12) = 6 + (–12) A. –18 B. –6 C. 6 D. 18

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

oh ok thx

OpenStudy (anonymous):

you welcome